Bankruptcy Risk and Z-Score Analysis
Following are selected ratios for JetBlue Airways for two recent fiscal years.

Ratio 2016 2015
Current ratio 0.604 0.705
Working capital to total assets (0.069) (0.104)
Retained earnings to total assets 0.258 0.194
EBIT to total assets 0.138 0.141
EBITDA to total assets 0.180 0.181
Market value of equity to total liabilities 1.380 1.342
Sales to total assets 0.699 0.742

Compute and interpret Altman Z-scores for the company for both years. (Do not round until your final answer; then round your answers to two decimal places.)

2016 z-score =
2015 z-score =
Answer #1

Answer:

The Altman Z-Score Formula

The original formula was created for publicly traded manufacturing companies.

Z-Score = 1.2(A) + 1.4(B) + 3.3(C) + 0.6(D) + 1.0(E)

Where:

A = Working Capital (Current Assets – Current liabilities) / Total Assets (Measures liquidity of firm)

B= Retained Earnings / Total Assets (measures accumulated profits compared to assets)

C= Earnings Before Interest & Taxes (EBIT) / Total Assets (measures how much profit the firms assets are producing)

D= Market Value of Equity (Mkt. Cap. + Preferred Stock) / Total Liabilities (compares the company’s value versus it’s liabilities)

E= Sales / Total Assets (efficiency ratio – measures how much the company’s assets are producing in sales).


2016 z-score = 1.2(-0.069)+1.4(0.258)+3.3(0.138)+0.6(1.380)+1.0(0.699) = 2.1576 = 2.16 rounded off.

2015 z-score = 1.2(-0.104)+1.4(0.194)+3.3(0.141)+0.6(1.342)+1.0(0.742) = 2.1593 = 2.16 rounded off.

Z-Score Results:

Z-Score of < 1.81 represents a company in distress.

Z-Score between 1.81 and 2.99 represents the “caution” zone.

Z-Score of over 3.0 represents a company with a safe balance sheet.

Thank you.